OpenStudy (anonymous):

given u = (1,-1,3) and v =(-4,5,3) find an ordered triple that represents 3u -2v a) (16,1,-7) b) (11,-13,3) c) (10,-12,1) d) (7,-7,16)

OpenStudy (anonymous):

@surjithayer here it is mane

OpenStudy (anonymous):

how did u get that @SA-Michael

OpenStudy (anonymous):

3u-2v=(3,-3,9)-(-8,10,6)=(3+8,-3-10,9-6)=(11,-13,3)

OpenStudy (anonymous):

thanks guys :) @SA-Michael @surjithayer

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

3u-2v=3(1;-1;3) -2(-4;5;3) =(3;-3;9) - (-8;10;6) =(3+8;-3-10;9-6) =(11;-13;3) Too late huh? xD